# Deterministic Dynamic Maximal Matching in Sublinear Update Time

### v1

@@ 4.3 Maintaining a Maximal Matching in the Adjunct Subgraph
Lemma 4.7 follows from (2) and (3).
4.4 Putting Everything Together: How to Derive Theorem 1.1
Since $E(H_{\texttt{core}})\subseteq E$, it follows that $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$ is a valid matching in $G$ (see Lemma 4.5). Furthermore, Lemma 4.7 guarantees that $M_{\texttt{adj}}$ is a maximal matching in $G_{\texttt{adj}}$ Thus, by Observation 4.6, the matching $M_{\texttt{final}}:=M_{\texttt{base}}\cup M_{\texttt{adj}}$ is a maximal matching in $G$. It now remains to analyze the overall amortized update time of our algorithm.
Recall that maintaining the EDCS $H$ incurs an update time of $O\left(\frac{n}{\epsilon B}\right)=O\left(\frac{n}{\delta B}\right)$, according to (1). Now, for the deterministic algorithm, Lemma 4.5 incurs an amortized update time of $\tilde{O}\left(B\cdot n^{1/2}\cdot\delta^{-3/2}\right)$. Thus, by Lemma 4.7, the overall amortized update time becomes
@@
For the randomized algorithm against an adaptive adversary, Lemma 4.5 incurs an amortized update time of $\tilde{O}(B\delta^{-1}+\delta^{-3})$. Thus, by Lemma 4.7, the overall amortized update time becomes
@@
This leads us to Theorem 1.1 in the decremental setting.
4.5 Dealing with Medium Nodes: Getting Rid of Assumption 4.1
We now provide a high-level outline of our algorithm in the general case, when Assumption 4.1 does not hold. Recall that under Assumption 4.1 there cannot be any node $v$ with $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H_{\texttt{init}}}(v)\leq\left(\frac{1}{2}+\delta\right)B$. Conceptually, the most significant challenge here is to deal with the set of medium nodes, given by:
@@
For the rest of this overview, we replace Assumption 4.1 with the following weaker assumption
Assumption 4.11 (Replaces Assumption 4.1).
@@
We now show how to adapt our algorithm to also handle the medium nodes. Lemma 4.12, stated below, generalizes Lemma 4.5, and serves as a key building block of our algorithm. The only difference between these two lemmas is that in Item 1 of Lemma 4.12, we allow for $O(\delta n)$ medium nodes that are free under $M_{\texttt{base}}$. Note that sets $V_{\texttt{sf}}$ and $V_{\texttt{dmg}}$ are defined the same as before (Definition 4.3), and are in particular both subsets of $V_{\texttt{v-hi}}$.
@@
Before outlining the proof sketch of Lemma 4.12, we explain how to adapt the algorithm for maintaining a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in the adjunct graph $G_{\texttt{adj}}:=(V_{\texttt{adj}},E_{\texttt{adj}})$, which is the subgraph of $G$ induced by the node-set $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$ (see Section 4.3). Lemma 4.13, stated below, generalizes Lemma 4.7 in the presence of medium nodes.
@@
Suppose that we maintain the matching $M_{\texttt{base}}$ as per Lemma 4.12. Then with an additive overhead of $\tilde{O}\left(\delta n+\frac{B}{\delta}+\frac{n}{\delta B}\right)$ amortized update time, we can deterministically and explicitly maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$.
@@
The only difference with the proof of Lemma 4.7 is that now the set $V_{\texttt{adj}}$ can contain some medium nodes, although $|V_{\texttt{adj}}\cap V_{\texttt{med}}|=O(\delta n)$ as per Item 1 of Lemma 4.12.
@@
We explicitly maintain the set $V_{\texttt{adj}}\cap V_{\texttt{med}}$ as a doubly linked list. Whenever a node $v\in V_{\texttt{lo}}\cup V_{\texttt{dmg}}$ is searching for a free neighbor (in the proof of Lemma 4.7), we ensure that it additionally scans the set $V_{\texttt{adj}}\cap V_{\texttt{med}}$. This leads to an additive overhead of $|V_{\texttt{med}}\cap V_{\texttt{adj}}|=O(\delta n)$ in the update time.
@@
Next, recall that $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$. Consider any edge $(u,v)\in E_{\texttt{adj}}\subseteq E\subseteq E_{\texttt{init}}$ with $v\in V_{\texttt{med}}$ and $u\in V_{\texttt{lo}}$. Then, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B+\left(\frac{1}{2}+\delta-\epsilon\right)B=(1-\epsilon)B$, and hence Definition 2.1 implies that $(u,v)\in E(H_{\texttt{init}})$. In other words, all the edges in $G_{\texttt{adj}}$ that connect a low node with a medium node belong to the EDCS $H_{\texttt{init}}$, which in turn has maximum degree at most $B$ (see Definition 2.1). Thus, every medium node has at most $B$ low neighbors in $G_{\texttt{adj}}$.
@@
Now, suppose that we are searching for a free neighbor of a node $v\in V_{\texttt{med}}$ in $G_{\texttt{adj}}$. This step involves scanning through: (i) all the low neighbors of $v$ in $G_{\texttt{adj}}$ (and by the previous discussion there are at most $B$ such neighbors), (ii) the set $V_{\texttt{med}}\cap V_{\texttt{adj}}$ of size $O(\delta n)$, and (iii) and at most $O\left(\frac{n}{B}\right)$ damaged nodes $V_{\texttt{dmg}}\cap V_{\texttt{adj}}$ (see Section 4.3). Overall, this scan takes $O\left(\delta n+B+\frac{n}{B}\right)$ time, and just as before, contributes an additive overhead of $O(\delta n)$ to the update time in Lemma 4.7.
Everything else remains the same as in the proof of Lemma 4.7. ∎
@@
Comparing Lemmas 4.12 and 4.13 against Lemmas 4.5 and 4.7, we conclude that there is an additive overhead of $O(\delta n)$ update time while dealing with the medium nodes. Finally, from the analysis from Section 4.4, it is easy to verify that this overhead of $O(\delta n)$ does not degrade our overall asymptotic update time (the values of $\delta,B$ remain the same as in Section 4.4).
Proof Sketch of Lemma 4.12
@@
Consider the scenario at the start of a new phase (just before the first update within that phase). At this point, we first compute a subgraph $H^{\prime}_{\texttt{init}}:=(V,E(H^{\prime}_{\texttt{init}}))$ of $H_{\texttt{init}}$, as follows. We initialize $H^{\prime}_{\texttt{init}}\leftarrow H_{\texttt{init}}$. Subsequently, for every node $v\in V_{\texttt{v-hi}}$, we keep deleting edges $(u,v)\in E(H^{\prime}_{\texttt{init}})$ incident on $v$ from $H^{\prime}_{\texttt{init}}$, until $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)$ becomes equal to $\Delta:=\left(\frac{1}{2}+\delta\right)B$. By Item 1 of Proposition 4.2, every edge $(u,v)$ that gets deleted from $H^{\prime}_{\texttt{init}}$ has its other endpoint $u\in V_{\texttt{lo}}$. Accordingly, this process does not reduce the degree of any medium node. When this for loop terminates, we have $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)\leq\Delta$ for all $v\in V_{\texttt{med}}\cup V_{\texttt{v-hi}}$, and $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$ for all $v\in V_{\texttt{lo}}$. In other words, the degree of every node $v\in V_{\texttt{med}}\cup V_{\texttt{v-hi}}$ in $H^{\prime}_{\texttt{init}}$ is at least $(1-\Theta(\delta))$ times the maximum degree $\Delta$. Now, we apply Lemma 3.1, with $\kappa=\Theta(\delta)$, to compute a matching $M\subseteq E(H^{\prime}_{\texttt{init}})\subseteq E(H_{\texttt{init}})$ that matches all but $O(\delta n)$ nodes in $V_{\texttt{med}}\cup V_{\texttt{v-hi}}$. Note that $\Delta=\Theta(B)$, and that a phase lasts for $\delta n$ updates in $G$. Thus, at the start of a phase, computing the subgraph $H^{\prime}_{\texttt{init}}$ and the invocation of Lemma 3.1 incurs an amortized update time of
@@
if we use the randomized algorithm. It is easy to verify that both these terms get subsumed within the respective (deterministic or randomized) guarantees of Lemma 4.12.
We have thus shown to initialize the desired matching $M_{\texttt{base}}=M$ at the beginning of the phase. We now give a sketch of how we maintain $M_{\texttt{base}}$ dynamically throughout the phase. Define $H_{\texttt{hilo}}$ to be the induced graph $H_{\texttt{core}}[V_{\texttt{v-hi}}\cup V_{\texttt{lo}}]$. Recall that all edges in $E_{H_{\texttt{core}}}(V_{\texttt{v-hi}})$ have their other endpoint in $V_{\texttt{lo}}$ (Item 1 of Proposition 4.2), so the set of edges incident to $V_{\texttt{v-hi}}$ is exactly the same in $E_{H_{\texttt{hilo}}}$ and $E_{H_{\texttt{core}}}$. We are thus able to apply Lemma 4.5 exactly as before to maintain a matching $M_{\texttt{hilo}}$ in $H_{\texttt{hilo}}$ that matches all of $V_{\texttt{sf}}$. We will maintain the invariant that $M_{\texttt{hilo}}\subseteq M_{\texttt{base}}$, thus guaranteeing that all nodes in $V_{\texttt{sf}}$ remain matched throughout the phase. The only remaining concern is that the algorithm for maintaining $M_{\texttt{hilo}}$ might end up unmatching medium vertices in $M_{\texttt{base}}$: if vertex $u\in V_{\texttt{hi}}$ becomes unmatched, then to maintain $M_{\texttt{hilo}}$ the algorithm from Lemma 4.5 finds an augmenting path from $u$ to some $v\in V_{\texttt{lo}}$ that was free in $M_{\texttt{hilo}}$; but although $v$ was free in $M_{\texttt{hilo}}$, it might have been matched to some medium node $x$ in $M_{\texttt{base}}$, in which case node $x$ now becomes unmatched. Fortunately, it is not hard to show that every adversarial update creates only $O(1)$ free vertices, so since $M_{\texttt{base}}$ has $O(\delta n)$ free medium vertices at the beginning of the phase, and since each phase lasts for $O(\delta n)$ updates, we maintain the desired property that there are always $O(\delta n)$ free medium vertices.
4.6 Matching All the Safe Nodes: Proof Outline for Lemma 4.5
@@
Lemma 4.14 (Simplified version of Lemma 5.1).
Let $G=(L\cup R,E)$ be a bipartite graph where, for some $X>0$ and $10<\gamma<$ 1, every vertex $v\in L$ has $2X\geq\mathrm{deg}(v)\geq X$ and every vertex in $R$ has degree $\leq X(1-\gamma)$. Suppose that the above degree constraints always hold as $G$ is subject to $O(\delta n)$ decremental updates. Then, there exists an algorithm that maintains a left-perfect matching in $G$ – i.e. a matching in which every vertex in $L$ is matched – with the following update times:
@@
Before proving the above Lemma, let us see why it implies Lemma 4.5
Proof sketch of Lemma 4.5.
For this proof sketch, we assume that we always have $V_{\texttt{v-hi}}=V_{\texttt{sf}}$; in the full version, we are able to effectively ignore damaged nodes because Lemma 4.5 only requires us to match all the nodes of $V_{\texttt{sf}}$. Let $H_{\texttt{hilo}}$ be the subgraph of $H_{\texttt{core}}$ that contains only the edges incident to $V_{\texttt{v-hi}}$. Observe that under our assumption of $V_{\texttt{v-hi}}=V_{\texttt{sf}}$, $H_{\texttt{hilo}}$ satisfies the properties of Lemma 4.14 with $X=B/2$ and $\gamma\sim\delta$: note in particular that the first property of an EDCS (Definition 2.1) ensures that every vertex has degree $\leq B=2X$, and that $H_{\texttt{hilo}}$ is bipartite because by Proposition 4.2, $H_{\texttt{init}}$ contains no edges between vertices in $V_{\texttt{v-hi}}$. We can thus apply Lemma 4.14 to maintain a matching that matches all of $V_{\texttt{sf}}$, as desired. The update-time bounds of Lemma 4.14 precisely imply those of Lemma 4.5; for the deterministic algorithm, recall that we amortize over the $O(\delta n)$ updates of a phase.
4.6.1 Maintaining Left-Perfect Matching: Proof of Lemma 4.14
@@
The deterministic algorithm is more involved than the randomized one, so we only give a high-level sketch. We start by maintaining a standard directed residual graph $G_{\textrm{res}}$ corresponding to the matching $M$ that we maintain (edges in $E\setminus M$ point from $L$ to $R$ in $G_{\textrm{res}}$, while edges in $M$ point from $R$ to $L$). We also create a dummy sink $t$ in $G_{\textrm{res}}$, with an edge from all $M$-free vertices in $R$ to $t$. It is not hard to check that there is a correspondence between paths from $L$ to $t$ in $G_{\textrm{res}}$ and augmenting paths in $G$. Our algorithm maintains a shortest path tree $T_{\mathrm{res}}$ to sink $t$ in $G_{\textrm{res}}$. Given any $M$-free free vertex $v\in L$, we can find an augmenting path from $v$ by simply following the $v-t$ path in $T_{\mathrm{res}}$.
@@
There exists a classic data structure known as an Even and Shiloach tree (denoted ES-tree) that can maintain a low-depth shortest path tree to a sink $t$ in a decremental graph [SE81]. The ES-tree can also handle a special kind of edge insertion into $G$ – namely, an edge insertion is a valid update if it does not decrease any distances to $t$. If $G_{\textrm{res}}$ were a decremental graph, then we could us an ES-tree to maintain $T_{\mathrm{res}}$, and we would be done. Unfortunately, even though the input graph $G$ to Lemma 4.14 only undergoes edge deletions, the graph $G_{\textrm{res}}$ can undergo edge insertions for two reasons.

### v2

@@ 4.3 Maintaining a Maximal Matching in the Adjunct Subgraph
Lemma˜4.7 follows from (2) and (3).
4.4 Putting Everything Together: How to Derive Theorem˜1.1
Since $E(H_{\texttt{core}})\subseteq E$, it follows that $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$ is a valid matching in $G$ (see Lemma˜4.5). Furthermore, Lemma˜4.7 guarantees that $M_{\texttt{adj}}$ is a maximal matching in $G_{\texttt{adj}}$ Thus, by Observation˜4.6, the matching $M_{\texttt{final}}:=M_{\texttt{base}}\cup M_{\texttt{adj}}$ is a maximal matching in $G$. It now remains to analyze the overall amortized update time of our algorithm.
Recall that maintaining the EDCS $H$ incurs an update time of $O\left(\frac{n}{\epsilon B}\right)=O\left(\frac{n}{\delta B}\right)$, according to (1). Now, for the deterministic algorithm, Lemma˜4.5 incurs an amortized update time of $\tilde{O}\left(B\cdot n^{1/2}\cdot\delta^{-3/2}\right)$. Thus, by Lemma˜4.7, the overall amortized update time becomes
@@
For the randomized algorithm against an adaptive adversary, Lemma˜4.5 incurs an amortized update time of $\tilde{O}(B\delta^{-1}+\delta^{-3})$. Thus, by Lemma˜4.7, the overall amortized update time becomes
@@
This leads us to Theorem˜1.1 in the decremental setting.
4.5 Dealing with Medium Nodes: Getting Rid of Assumption˜4.1
We now provide a high-level outline of our algorithm in the general case, when Assumption˜4.1 does not hold. Recall that under Assumption˜4.1 there cannot be any node $v$ with $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H_{\texttt{init}}}(v)\leq\left(\frac{1}{2}+\delta\right)B$. Conceptually, the most significant challenge here is to deal with the set of medium nodes, given by:
@@
For the rest of this overview, we replace Assumption˜4.1 with the following weaker assumption
Assumption 4.11 (Replaces Assumption˜4.1).
@@
We now show how to adapt our algorithm to also handle the medium nodes. Lemma˜4.12, stated below, generalizes Lemma˜4.5, and serves as a key building block of our algorithm. The only difference between these two lemmas is that in Item˜1 of Lemma˜4.12, we allow for $O(\delta n)$ medium nodes that are free under $M_{\texttt{base}}$. Note that sets $V_{\texttt{sf}}$ and $V_{\texttt{dmg}}$ are defined the same as before (Definition˜4.3), and are in particular both subsets of $V_{\texttt{v-hi}}$.
@@
Before outlining the proof sketch of Lemma˜4.12, we explain how to adapt the algorithm for maintaining a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in the adjunct graph $G_{\texttt{adj}}:=(V_{\texttt{adj}},E_{\texttt{adj}})$, which is the subgraph of $G$ induced by the node-set $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$ (see Section˜4.3). Lemma˜4.13, stated below, generalizes Lemma˜4.7 in the presence of medium nodes.
@@
Suppose that we maintain the matching $M_{\texttt{base}}$ as per Lemma˜4.12. Then with an additive overhead of $\tilde{O}\left(\delta n+\frac{B}{\delta}+\frac{n}{\delta B}\right)$ amortized update time, we can deterministically and explicitly maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$.
@@
The only difference with the proof of Lemma˜4.7 is that now the set $V_{\texttt{adj}}$ can contain some medium nodes, although $|V_{\texttt{adj}}\cap V_{\texttt{med}}|=O(\delta n)$ as per Item˜1 of Lemma˜4.12.
@@
We explicitly maintain the set $V_{\texttt{adj}}\cap V_{\texttt{med}}$ as a doubly linked list. Whenever a node $v\in V_{\texttt{lo}}\cup V_{\texttt{dmg}}$ is searching for a free neighbor (in the proof of Lemma˜4.7), we ensure that it additionally scans the set $V_{\texttt{adj}}\cap V_{\texttt{med}}$. This leads to an additive overhead of $|V_{\texttt{med}}\cap V_{\texttt{adj}}|=O(\delta n)$ in the update time.
@@
Next, recall that $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$. Consider any edge $(u,v)\in E_{\texttt{adj}}\subseteq E\subseteq E_{\texttt{init}}$ with $v\in V_{\texttt{med}}$ and $u\in V_{\texttt{lo}}$. Then, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B+\left(\frac{1}{2}+\delta-\epsilon\right)B=(1-\epsilon)B$, and hence Definition˜2.1 implies that $(u,v)\in E(H_{\texttt{init}})$. In other words, all the edges in $G_{\texttt{adj}}$ that connect a low node with a medium node belong to the EDCS $H_{\texttt{init}}$, which in turn has maximum degree at most $B$ (see Definition˜2.1). Thus, every medium node has at most $B$ low neighbors in $G_{\texttt{adj}}$.
@@
Now, suppose that we are searching for a free neighbor of a node $v\in V_{\texttt{med}}$ in $G_{\texttt{adj}}$. This step involves scanning through: (i) all the low neighbors of $v$ in $G_{\texttt{adj}}$ (and by the previous discussion there are at most $B$ such neighbors), (ii) the set $V_{\texttt{med}}\cap V_{\texttt{adj}}$ of size $O(\delta n)$, and (iii) and at most $O\left(\frac{n}{B}\right)$ damaged nodes $V_{\texttt{dmg}}\cap V_{\texttt{adj}}$ (see Section˜4.3). Overall, this scan takes $O\left(\delta n+B+\frac{n}{B}\right)$ time, and just as before, contributes an additive overhead of $O(\delta n)$ to the update time in Lemma˜4.7.
Everything else remains the same as in the proof of Lemma˜4.7. ∎
@@
Comparing Lemmas˜4.12 and 4.13 against Lemmas˜4.5 and 4.7, we conclude that there is an additive overhead of $O(\delta n)$ update time while dealing with the medium nodes. Finally, from the analysis from Section˜4.4, it is easy to verify that this overhead of $O(\delta n)$ does not degrade our overall asymptotic update time (the values of $\delta,B$ remain the same as in Section˜4.4).
Proof Sketch of Lemma˜4.12
@@
Consider the scenario at the start of a new phase (just before the first update within that phase). At this point, we first compute a subgraph $H^{\prime}_{\texttt{init}}:=(V,E(H^{\prime}_{\texttt{init}}))$ of $H_{\texttt{init}}$, as follows. We initialize $H^{\prime}_{\texttt{init}}\leftarrow H_{\texttt{init}}$. Subsequently, for every node $v\in V_{\texttt{v-hi}}$, we keep deleting edges $(u,v)\in E(H^{\prime}_{\texttt{init}})$ incident on $v$ from $H^{\prime}_{\texttt{init}}$, until $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)$ becomes equal to $\Delta:=\left(\frac{1}{2}+\delta\right)B$. By Item˜1 of Proposition˜4.2, every edge $(u,v)$ that gets deleted from $H^{\prime}_{\texttt{init}}$ has its other endpoint $u\in V_{\texttt{lo}}$. Accordingly, this process does not reduce the degree of any medium node. When this for loop terminates, we have $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)\leq\Delta$ for all $v\in V_{\texttt{med}}\cup V_{\texttt{v-hi}}$, and $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$ for all $v\in V_{\texttt{lo}}$. In other words, the degree of every node $v\in V_{\texttt{med}}\cup V_{\texttt{v-hi}}$ in $H^{\prime}_{\texttt{init}}$ is at least $(1-\Theta(\delta))$ times the maximum degree $\Delta$. Now, we apply Lemma˜3.1, with $\kappa=\Theta(\delta)$, to compute a matching $M\subseteq E(H^{\prime}_{\texttt{init}})\subseteq E(H_{\texttt{init}})$ that matches all but $O(\delta n)$ nodes in $V_{\texttt{med}}\cup V_{\texttt{v-hi}}$. Note that $\Delta=\Theta(B)$, and that a phase lasts for $\delta n$ updates in $G$. Thus, at the start of a phase, computing the subgraph $H^{\prime}_{\texttt{init}}$ and the invocation of Lemma˜3.1 incurs an amortized update time of
@@
if we use the randomized algorithm. It is easy to verify that both these terms get subsumed within the respective (deterministic or randomized) guarantees of Lemma˜4.12.
We have thus shown to initialize the desired matching $M_{\texttt{base}}=M$ at the beginning of the phase. We now give a sketch of how we maintain $M_{\texttt{base}}$ dynamically throughout the phase. Define $H_{\texttt{hilo}}$ to be the induced graph $H_{\texttt{core}}[V_{\texttt{v-hi}}\cup V_{\texttt{lo}}]$. Recall that all edges in $E_{H_{\texttt{core}}}(V_{\texttt{v-hi}})$ have their other endpoint in $V_{\texttt{lo}}$ (Item˜1 of Proposition˜4.2), so the set of edges incident to $V_{\texttt{v-hi}}$ is exactly the same in $E_{H_{\texttt{hilo}}}$ and $E_{H_{\texttt{core}}}$. We are thus able to apply Lemma˜4.5 exactly as before to maintain a matching $M_{\texttt{hilo}}$ in $H_{\texttt{hilo}}$ that matches all of $V_{\texttt{sf}}$. We will maintain the invariant that $M_{\texttt{hilo}}\subseteq M_{\texttt{base}}$, thus guaranteeing that all nodes in $V_{\texttt{sf}}$ remain matched throughout the phase. The only remaining concern is that the algorithm for maintaining $M_{\texttt{hilo}}$ might end up unmatching medium vertices in $M_{\texttt{base}}$: if vertex $u\in V_{\texttt{hi}}$ becomes unmatched, then to maintain $M_{\texttt{hilo}}$ the algorithm from Lemma˜4.5 finds an augmenting path from $u$ to some $v\in V_{\texttt{lo}}$ that was free in $M_{\texttt{hilo}}$; but although $v$ was free in $M_{\texttt{hilo}}$, it might have been matched to some medium node $x$ in $M_{\texttt{base}}$, in which case node $x$ now becomes unmatched. Fortunately, it is not hard to show that every adversarial update creates only $O(1)$ free vertices, so since $M_{\texttt{base}}$ has $O(\delta n)$ free medium vertices at the beginning of the phase, and since each phase lasts for $O(\delta n)$ updates, we maintain the desired property that there are always $O(\delta n)$ free medium vertices.
4.6 Matching All the Safe Nodes: Proof Outline for Lemma˜4.5
@@
Lemma 4.14 (Simplified version of Lemma˜5.1).
Let $G=(L\cup R,E)$ be a bipartite graph where, for some $X>0$ and $0<\gamma<1$, every vertex $v\in L$ has $2X\geq\mathrm{deg}(v)\geq X$ and every vertex in $R$ has degree $\leq X(1-\gamma)$. Suppose that the above degree constraints always hold as $G$ is subject to $O(\delta n)$ decremental updates. Then, there exists an algorithm that maintains a left-perfect matching in $G$ – i.e. a matching in which every vertex in $L$ is matched – with the following update times:
@@
Before proving the above Lemma, let us see why it implies Lemma˜4.5
Proof sketch of Lemma˜4.5.
For this proof sketch, we assume that we always have $V_{\texttt{v-hi}}=V_{\texttt{sf}}$; in the full version, we are able to effectively ignore damaged nodes because Lemma˜4.5 only requires us to match all the nodes of $V_{\texttt{sf}}$. Let $H_{\texttt{hilo}}$ be the subgraph of $H_{\texttt{core}}$ that contains only the edges incident to $V_{\texttt{v-hi}}$. Observe that under our assumption of $V_{\texttt{v-hi}}=V_{\texttt{sf}}$, $H_{\texttt{hilo}}$ satisfies the properties of Lemma˜4.14 with $X=B/2$ and $\gamma\sim\delta$: note in particular that the first property of an EDCS (Definition˜2.1) ensures that every vertex has degree $\leq B=2X$, and that $H_{\texttt{hilo}}$ is bipartite because by Proposition˜4.2, $H_{\texttt{init}}$ contains no edges between vertices in $V_{\texttt{v-hi}}$. We can thus apply Lemma˜4.14 to maintain a matching that matches all of $V_{\texttt{sf}}$, as desired. The update-time bounds of Lemma˜4.14 precisely imply those of Lemma˜4.5; for the deterministic algorithm, recall that we amortize over the $O(\delta n)$ updates of a phase.
4.6.1 Maintaining Left-Perfect Matching: Proof of Lemma˜4.14
@@
The deterministic algorithm is more involved than the randomized one, so we only give a high-level sketch. We start by maintaining a standard directed residual graph $G_{\textrm{res}}$ corresponding to the matching $M$ that we maintain (edges in $E\setminus M$ point from $L$ to $R$ in $G_{\textrm{res}}$, while edges in $M$ point from $R$ to $L$). We also create a dummy sink $t$ in $G_{\textrm{res}}$, with an edge from all $M$-free vertices in $R$ to $t$. It is not hard to check that there is a correspondence between paths from $L$ to $t$ in $G_{\textrm{res}}$ and augmenting paths in $G$. Our algorithm maintains a shortest path tree $T_{\mathrm{res}}$ to sink $t$ in $G_{\textrm{res}}$. Given any $M$-free free vertex $v\in L$, we can find an augmenting path from $v$ by simply following the vtv-t path in $T_{\mathrm{res}}$.
@@
There exists a classic data structure known as an Even and Shiloach tree (denoted ES-tree) that can maintain a low-depth shortest path tree to a sink $t$ in a decremental graph [SE81]. The ES-tree can also handle a special kind of edge insertion into $G$ – namely, an edge insertion is a valid update if it does not decrease any distances to $t$. If $G_{\textrm{res}}$ were a decremental graph, then we could us an ES-tree to maintain $T_{\mathrm{res}}$, and we would be done. Unfortunately, even though the input graph $G$ to Lemma˜4.14 only undergoes edge deletions, the graph $G_{\textrm{res}}$ can undergo edge insertions for two reasons.